# Mind the Error! Detection and Localization of Instruction Errors in Vision-and-Language Navigation

### v1

@@ VII Conclusion
We presented the novel benchmark R2RIE-CE, a new benchmark for VLN-CE, in which different types of errors are injected into textual instructions. Based on our experiments, state-of-the-art VLN-CE methods are affected by instruction perturbations. On top of R2RIE-CE, we also established the novel task of Detection and Localization of Instruction Errors. Our proposed method, IEDL, composed of detection and localization heads, can detect and localize errors within a sub-sentence distance on the original instruction. As a further experiment, by running IEDL on the standard R2R-CE and RxR-CE, we were able to find $8$ and $10$ episodes, respectively, that should be removed from the validation set. Our benchmark and proposed method support the development of more reliable and versatile agents capable of handling real-world ambiguity and uncertainty. In future work, we will investigate error-aware policy learning to improve the navigation performance in the VLN-CE task.

### v2

@@ VII Conclusion
We presented the novel benchmark R2RIE-CE, a new benchmark for VLN-CE, in which different types of errors are injected into textual instructions. Based on our experiments, state-of-the-art VLN-CE methods are affected by instruction perturbations. On top of R2RIE-CE, we also established the novel task of Detection and Localization of Instruction Errors. Our proposed method, IEDL, composed of detection and localization heads, can detect and localize errors within a sub-sentence distance on the original instruction. As a further experiment, by running IEDL on the standard R2R-CE and RxR-CE, we were able to find $8$ and $10$ episodes, respectively, that should be removed from the validation set. Our benchmark and proposed method support the development of more reliable and versatile agents capable of handling real-world ambiguity and uncertainty. In future work, we will investigate error-aware policy learning to improve the navigation performance in the VLN-CE task, and online instructions errors detection [31].